2 COSTS AND COSTING SYSTEMS 5.26 Weighted average cost and FIFO: manufacturer The following da During Februar 89000nits were completed and transferred out to Australian Packaging Ltd, a manufacturer of cardboard boxes: LO31 10000 units 10900 28950 Work in process, 1 February: Direct material Conversion Costs incurred during February: Direct material Conversion $112700 160200 complete as to direct material: 35% complete as to conversion. The equivalent units for February were as follows: Weighted average FIFO 93000 93500 103000 Direct material Conversion | 97000-lo.メ0 Required: 1 Calculate each of the following amount ed e pro g: (a) cost of goods completed during February. (b) cost of the 28 February work in process inventory. Repeat requirement 1 using the FIFO method. 2
Answer #1

1)Using Weighted Average Method

Cost Per Equivalent Unit (Direct Material) =Total Material Cost/ Equivalent Unit of Direct material

=(10,900+112,700)/103,000

=1.20

Cost Per Equivalent Unit (Conversion) =Total Conversion Cost/ Equivalent Unit of Conversion

=(28,950+160,200)/97,000

=1.95

Equivalent unit of Direct Material in Ending Work in process = Total Equivalent Unit of Direct Material - Unit completed & Transferred out

=103,000-89,000 =14,000

Equivalent unit of Conversion in Ending Work in process = Total Equivalent Unit of Conversion - Unit completed & Transferred out

=97,000 -89,000 =8,000

Cost of Goods Completed and transferred out = Unit completed & Transferred out*(Cost per Equivalent Unit of Direct Material + Cost per Equivalent Unit of Conversion)

=89,000*(1.20+1.95)

=280,350

Cost of Feb 28 Work in Process Inventory= Equivalent unit of Direct Material in Ending Work in process *Cost per Equivalent Unit of Direct Material + Equivalent unit of Conversion in Ending Work in process * Cost per Equivalent Unit of Conversion

=(14,000 x 1.20) +(8,000 x 1.95)

=32,400

2)Using  FIFO METHOD

Cost per Equivalent Unit -Direct Material = Material cost in Feb/Equivalent Unit of Direct Material

=112,700/93000

1.2118288

Cost per Equivalent Unit -Conversion = Conversion cost in Feb/Equivalent Unit of Conversion

=160,200/93,500

=1.713369

Equivalent unit of Direct Material in Ending Work in process would remain same as weighted average method i.e 14000 Units

Similarly

Equivalent unit of Conversion in Ending Work in process would remain same as weighted average method i.e 8000 Units

Unit Started and completed = Unit Completed & transferred out - Beginning Work in process

=89,000-10,000

=79,000

Equivalent unit of Direct Material in Beginning Work in process during Feb= Total Equivalent Unit of Direct Material- Equivalent unit of Direct Material in Ending Work in process- Unit Started and completed

=93,000-14,000-79,000 =0

Equivalent unit of Conversion in Beginning Work in process during Feb= Total Equivalent Unit of Conversion- Equivalent unit of Conversion in Ending Work in process- Unit Started and completed

=93,500-8,000-79,000

=6,500

Cost of Goods Completed and transferred out = Beginning Work in Process + Cost incurred to complete BWIP + Unit Started & completed*(Cost per Equivalent Unit of Direct Material + Cost per Equivalent Unit of Conversion)

=(10,900+28,950) +(0x 1.211828+6,500x 1.713369)+ 79,000 x(1.211828+1.713369)

=282,077.50

Cost of Feb 28 Work in Process Inventory= Equivalent unit of Direct Material in Ending Work in process *Cost per Equivalent Unit of Direct Material + Equivalent unit of Conversion in Ending Work in process * Cost per Equivalent Unit of Conversion

14,000 x 1.211828+ 8,000 x 1.713369

=30,672.54